OpenStudy (anonymous):

A line has equation y+4=2 what is the slope of the line? What could be the coordinates of two points on the line?

OpenStudy (jadedry):

Is that y+4 = 2x or just y+4 = 2?

OpenStudy (jadedry):

or even y+4x = 2 ?

OpenStudy (jadedry):

Because if it is y +4 = 2 then y = - 2 Therefore the line is parallel to the x axis. (It does not intercept it)

OpenStudy (jadedry):

In which case the slope would be zero, and any point on the line would be (n,-2) where you substitute n with a number of your choice.